# Automated Privacy Information Annotation in Large Language Model Interactions

### v1

@@ 4.2 Pipeline Design
Extensive Privacy Categories Extraction. We perform basic extraction for each sample labeled as leaking privacy. We prompt GPT-4 to directly identify privacy phrases and the corresponding categories. We also provide an example to guide GPT-4. These results are then merged to create an extensive set of privacy categories. However, since extraction for each sample is independent, some categories may be expressed differently but hold the same meaning. For example, we may encounter both “hobby” and “interests and hobbies,” which should be treated as equivalent. To address this, we use GPT-4 to identify such duplicates and assist in the merging process.
@@
Specifically, as shown in Figure 3, we partition the shuffled privacy categories into smaller, more manageable blocks, and deduplicate categories within each block independently. The deduplicated results from all blocks are then merged. This process is repeated iteratively until no further deduplication occurs, meaning the number of categories remains unchanged before and after the deduplication step. To reduce randomness, we run the algorithm twice and perform a final deduplication by prompting GPT-4 to deduplicate across all categories at once. In practice, we find that using only one-tenth of the data for category extraction yields results that are nearly saturated in terms of diversity.
@@
Privacy Phrase Extraction. For each sample that contains privacy leakage, based on the expensive set of privacy categories, we prompt GPT-4 to extract phrases that explicitly reveal privacy information about the user. By inserting categories into the prompts, we ask GPT-4 to match the phrases and the categories. The extracted phrase must be directly copied from the query, and the relationship between the phrase and the user should be evident, not speculative. For example, phrases should not be extracted with the reason like “the user may be interested in something.” Additionally, any phrases falling outside the given privacy categories, which may indicate an wrong extraction, should be excluded. To enhance the precision of the extracted phrases, we apply two filtering rules: (1) the privacy phrase must be directly associated with the user or their close associates, and (2) the phrase must have an explicit reference, rather than be a general phrase such as “go a place” or “have ideas”. We use GPT-4 to evaluate these criteria and remove any phrases that do not meet the rules. The results of experiments on filtering using these rules are shown in Table 1.
@@
In practice, samples misclassified as leaking privacy will be corrected if no phrases are extracted. Furthermore, including too many categories in a single input can reduce both the precision and the recall of phrases. To mitigate this, we divide the categories into smaller blocks and make multiple requests to GPT-4 for each sample, then merge the results. This approach may result in the extraction of synonymous phrases across different requests, such as “plan to play football” and “play football.” Therefore, we prompt GPT-4 to deduplicate the results, retaining the more concise and clear phrase.
@@
Privacy Information Annotation. For each extracted phrase, we prompt GPT-4 to annotate the specific privacy information it reveals, based on the context of whole text. For example, as shown in Figure 1, the privacy information associated with the phrase “London” is “The user lived in London when he was a child.” We also provide several manually constructed examples for annotation.
4.3 Pipeline over Dialogue Datasets
@@
Based on the automated annotation pipeline, we create a large dataset comprising over 249K queries from 33K dialogues with 154K privacy phrases and corresponding information in different languages.
@@
Raw Datasets. We use a diverse set of source corpora, including one English dialogue dataset, ShareGPT sharegpt , and three Chinese dialogue datasets, CrossWOZ crosswoz , DuConv duconv , and LCCC-base lccc . The large scale and varied scope of these datasets make them well-suited as our data source. Specifically, ShareGPT naturally consists of human-to-machine interactions, while the tree Chinese datasets consist of human-to-human dialogues. To adapt the human-to-human dialogues for our task, we designate the first speaker as the target user for privacy information extraction and the second speaker as the assistant. Detailed introduction of source corpora is provided in Appendix B.
Figure 4: Top 20 privacy categories in ShareGPT.
@@
Annotated Datasets. After each step of the automated annotation pipeline, we have: (1) classifying privacy leakage obtains over 41K samples with privacy leakage and 56K without privacy leakage in the English portion of dataset, and 64K samples with privacy leakage and 88K without privacy leakage in the Chinese portion; (2) extracting privacy categories from 2,000 English samples and 5,000 Chinese samples results in 325 categories for English and 149 for Chinese; (3) extracting privacy phrases and annotating information obtain 97,695 English samples and 151,988 Chinese samples, of which 32,814 English samples and 43,255 Chinese samples contain privacy, corresponding to 85,320 privacy phrases and information in English and 68,910 in Chinese. Appendix C shows a data example and the usage details of LLM API are presented in Appendix L.
Figure 5: Word cloud of annotated privacy information.
Dataset Analysis. Taking the English dataset as a case study, Figure 4 shows the relative frequency of top 20 different privacy categories. We can find that in open-domain interaction dialogues, information such as users’ status, opinions, preferences, and other potentially sensitive details is more likely to be inadvertently revealed, whereas categories like names appear less frequently.
Figure 5 presents a word cloud of the identified privacy information in English dataset. We can observe that the word cloud reveals a high frequency of terms associated with the user’s “involvement” in various activities, as well as their personal interests. These results indicates that the privacy information in user interactions is significantly different from PII in specific documents or content on the web.
Comparison with Human Annotation. To evaluate the privacy leakage classification step, we have 8 annotators classify 200 samples from ShareGPT, labeling them as either leak or non-leak. The detailed guidelines to ensure the objectivity and consistency of the annotation results are provided in Appendix F. The results show an accuracy of 0.95, a recall score of 0.94, and a precision score of 0.97. These metrics demonstrate that results of the first step closely align with the manual annotations.
@@
Table 1: Human evaluation results of phrase extraction, prompting GPT-4 vs. our pipeline.
[…]
Directly prompting GPT-4 68.08 90.99 77.89 +00.00
Phrase extraction in pipeline 98.67 82.90 90.10 +12.21
      + Filtering rules 96.30 90.17 93.13 +15.24
@@
We also conduct a human evaluation for the privacy phrase extraction step, as shown in Table 1. Specifically, we manually extract privacy phrases for 100 samples to serve as the ground truth, resulting in 222 privacy phrases. Using the evaluation method described in Section 4.4, we found that the recall score for phrases directly extracted by GPT-4 with prompts ghost was only 68.08%, meaning that roughly one-third of privacy phrases were missed. In comparison to this baseline, our pipeline, by incorporating category information as additional input, shows a decrease in precision but significantly increases the number of extracted phrases, boosting the recall score by 30.59%. As a result, the F1 score improves by 12.21%. Furthermore, applying filtering rules further enhances the F1 score by an additional 3.03%. These results suggest that the phrase extraction step aligns well with manual annotations, highlighting its effectiveness and consistency.
4.4 Evaluation Metrics
Privacy Query-Level Metric for Leakage Classification Task. Given an user’s query, the objective of privacy leakage classification task is to determine whether the query reveals any privacy of the user or user-related individuals. The automatically annotated data can be divided into either “leaked” or “non-leaked,” based on whether there are privacy phrases extracted. For such query-level classification task, we adopt accuracy to measure the performance of different local privacy detection methods.
Privacy Phrase-Level Metrics for Phrase Extraction Task. Given a private query, the objective of privacy phrase extraction task is to extract phrases based on the context of query that results in user’s privacy leakage. For each data sample, we focus on both missed and incorrect extractions. For missed extractions, we define the recall score of sample $q$ as the ratio of the intersection between the predicted phrases and the ground truth, normalized by the total number of ground truth phrases. For incorrect extractions, we define the precision score of $q$ as the ratio of the intersection between the predicted phrases and the ground truth, normalized by the total number of predicted phrases. The overall recall and precision scores are calculated as the average of these scores across all samples.
Specifically, we let $\mathcal{Q}_{r}$ denote the set of user queries where the extracted phrase is non-empty in the ground truth and $\mathcal{Q}_{p}$ denote the set of user queries where the extracted phrase is non-empty in the prediction results. We let $\hat{\mathcal{P}}_{q}$ and $\mathcal{P}_{q}$ denote the set of the predicted phrases and the ground truth of the query $q$, respectively. Then, the recall score $\text{R}_{\mathcal{P}}$ and precision score $\text{P}_{\mathcal{P}}$ is defined as
$\text{R}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\mathcal{P}_{q}|},\ \ \text{P}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{q\in\mathcal{Q}_{p}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\hat{\mathcal{P}}_{q}|},$ (1)
where ||$\cdot$|| ⋅ | is the size of the set. In practice, the extracted phrases may refer to the same entity but have different expressions. To address this issue, our phrase matching strategy goes beyond simply checking whether two phrases are identical. It also considers two phrases a match if one is part of the other or if their Rouge-L score exceeds 0.5 privacyimplications .
Privacy Information-Level Metrics for Information Summarization Task. Given a query with privacy phrase, the objective of privacy information summarization task is to summarize specific information that be leaked from the query. In general, we use Rouge-L to evaluate the quality of generated privacy information. Similarly, we define the miss and incorrect extractions of the privacy information. We let ^$\hat{\mathcal{I}}_{q}$ and $\mathcal{I}_{q}$ denote the set of the predicted information and the ground truth of the query $q$, respectively. For one privacy information ${I}_{j}\in{\mathcal{I}}_{q}$, the recall score is defined as $f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})=\max_{\hat{I}_{k}\in\hat{\mathcal{I}}_{q}}\text{{Rouge}}({I}_{j},\hat{I}_{k}),$ where, Rouge is the function that calculate the Rouge-L score for two sentences. The recall score of ${\mathcal{I}}_{q}$ is defined as the average recall score of each piece of information it contains. For $\hat{I}_{j}\in\hat{\mathcal{I}}_{q}$, the precision score is defined as $f_{\text{P}}(\hat{I}_{j},\mathcal{I}_{q})=\max_{I_{k}\in\mathcal{I}_{q}}\text{{Rouge}}(\hat{I}_{j},I_{k}).$ The precision score of ^$\hat{\mathcal{I}}_{q}$ is defined as the average precision score of each piece of information it contains. Then we can define the recall score $\text{R}_{\mathcal{I}}$ and the precision score $\text{P}_{\mathcal{I}}$ for all queries as follow,
$\text{R}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{1}{|\mathcal{I}_{D}|}{\sum_{I_{j}\in\mathcal{I}_{q}}f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})},\ \ \text{P}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{D\in\mathcal{Q}_{p}}\frac{1}{|\hat{\mathcal{I}}_{q}|}{\sum_{\hat{I}_{j}\in\hat{\mathcal{I}}_{q}}f_{\text{P}}(\hat{I}_{j},{\mathcal{I}}_{q})}.$ (2)
Intuitively, $\text{R}_{\mathcal{I}}$ measures the completeness of the extracted privacy information, while $\text{P}_{\mathcal{I}}$ evaluates the accuracy of the privacy information.
After getting the recall and precision score, we can also define the F1 score for both phrase-level and information-level as follow,
[…]
5 Evaluation for Local Privacy Detection Methods
Table 2: Accuracy of different baselines with query-level metric (%).
Method Model English Chinese
Acc Acc
ZG Llama-3.2-1B $61.18_{\pm 0.49}^{*}$ $58.08_{\pm 0.49}^{*}$
Llama-3.2-3B $65.07_{\pm 0.48}^{*}$ $66.97_{\pm 0.47}$
@@ 5 Evaluation for Local Privacy Detection Methods
Table 3: Performance of different privacy detection baselines with phrase-level metrics (%).
Method Model English Chinese
[…]
ZG Llama-3.2-1B $29.67_{\pm 0.43}^{*}$ $27.21_{\pm 0.38}^{*}$ 28.39 $22.07_{\pm 0.40}^{*}$ $26.08_{\pm 0.43}^{*}$ 23.91
@@
5.1 Setups

### v2

@@ 4.2. Pipeline Design
Extensive Privacy Categories Extraction. We first prompt GPT-4o to directly identify privacy phrases and the corresponding categories for each sample labeled as leaking privacy. We also provide an example to guide GPT-4o. These results are then merged to create an extensive set of privacy categories. However, since extraction for each sample is independent, some categories may be expressed differently but hold the same meaning. For example, we may encounter both “hobby” and “interests and hobbies,” which should be treated as equivalent. To address this, we use GPT-4o to identify such duplicates and assist in the merging process.
@@
Specifically, as shown in Figure 3, we partition the shuffled privacy categories into smaller, more manageable blocks and deduplicate categories within each block independently. The deduplicated results from all blocks are then merged. This process is repeated iteratively until no further deduplication occurs, meaning the number of categories remains unchanged before and after the deduplication step. To reduce randomness, we run the algorithm twice and perform a final deduplication by prompting GPT-4o to deduplicate across all categories at once. In practice, given a seen or even unseen dialogue dataset, performing fine-grained analysis on each sample and then merging ensures comprehensive category coverage at the dataset level and highlights the pipeline’s adaptability and generalizability across diverse dialogue domains. We also find that using only one-tenth of the data for category extraction yields results that are nearly saturated in terms of diversity.
@@
Privacy Phrase Extraction. For each sample that contains privacy leakage, based on the expensive set of privacy categories, we prompt GPT-4o to extract phrases that explicitly reveal privacy information about the user who poses the query. By inserting categories into the prompts, we ask GPT-4o to match the phrases and the categories. The extracted phrase must be directly copied from the query, and the relationship between the phrase and the user should be evident, not speculative. For example, phrases should not be extracted with the reason like “the user may be interested in something.” Additionally, any phrases falling outside the given privacy categories, which may indicate a wrong extraction, should be excluded. To enhance the precision of the extracted phrases, we apply two filtering rules: (1) the privacy phrase must be directly associated with the user or their close associates, and (2) the phrase must have an explicit reference, rather than be a general phrase such as “go a place” or “have ideas”. We use GPT-4o to evaluate these criteria and remove any phrases that do not meet the rules. The effectiveness of filtering based on the rules is shown in Table 2.
@@
In practice, samples misclassified as leaking privacy will be corrected if no phrases are extracted. Furthermore, including too many categories in a single input can reduce both precision and recall of phrases. To mitigate this, we divide the categories into smaller blocks and make multiple requests to GPT-4o for each sample, then merge the results. This approach may result in the extraction of synonymous phrases across different requests, such as “plan to play football” and “play football.” Therefore, we also prompt GPT-4o for deduplication, retaining the more concise and clearer phrase.
@@
Privacy Information Annotation. Finally, for each extracted phrase, we prompt GPT-4o to annotate the specific privacy information it reveals, based on the context of the whole text. We also provide several manually constructed examples for annotation.
4.3. Pipeline over Dialogue Datasets
@@
Based on the automated annotation pipeline, we create large datasets comprising over 249K queries from 33K dialogues with 154K privacy phrases and corresponding information in different languages.
@@
Raw Datasets. We use a diverse set of source corpora, including one English dialogue dataset, ShareGPT (ShareGPT, 2023), and three Chinese dialogue datasets, CrossWOZ (Zhu et al., 2020), DuConv (Wu et al., 2019), and LCCC-base (Wang et al., 2020). The large scale and varied scope of these datasets make them well-suited as our data source. Specifically, ShareGPT naturally consists of human-to-machine interactions, while the three Chinese datasets consist of human-to-human dialogues. To adapt the human-to-human dialogues for our task, we designate the first speaker as the target user for privacy information extraction and the second speaker as the assistant. A detailed introduction of the source corpora is provided in Appendix B.1.
(a) Top 20 privacy categories.
(b) Word cloud.
Figure 4. Analyses of privacy information in ShareGPT (ShareGPT, 2023).
(a) Top 20 privacy categories.
(b) Word cloud.
Figure 5. Analyses of privacy information in CrossWOZ (Zhu et al., 2020), DuConv (Wu et al., 2019), and LCCC-base (Wang et al., 2020).
@@
Annotated Datasets. After each step of the automated annotation pipeline, we have: (1) classifying privacy leakage obtains over 41K samples with privacy leakage and 56K without privacy leakage in the English portion of the dataset, and 64K samples with privacy leakage and 88K without privacy leakage in the Chinese portion; (2) extracting privacy categories from 2,000 English samples and 5,000 Chinese samples results in 325 categories for English and 149 for Chinese; (3) extracting privacy phrases and annotating information obtain 97,659 English samples and 151,988 Chinese samples, of which 32,814 English samples and 43,255 Chinese samples contain privacy, corresponding to 85,320 privacy phrases and information in English and 68,910 in Chinese. Appendix B.2 shows a data example, and the LLM API usage is presented in Appendix E.
Dataset Analysis. Figure 4(a) and Figure 5(a) show the relative frequency of the top 20 different privacy categories in the English and the Chinese datasets, respectively. We can find that in open-domain interactions, information such as users’ status, opinions, and preferences is more likely to be inadvertently revealed, whereas categories like age and gender appear less frequently. Figure 4(b) presents a word cloud of the identified privacy information in the English dataset. We can observe that the word cloud reveals a high frequency of terms associated with the user’s “involvement” in various activities, as well as their personal interests. These results indicate that the privacy information in user interactions with LLM is significantly different from PII in specific documents or content on the web. Additionally, a comparison of the word clouds across languages in Figure 4(b) and Figure 5(b) reveals scenario differences between the English and the Chinese datasets. While the English dataset predominantly features discussions centered on work or project-related topics, the Chinese datasets contain more casual, everyday conversations, often focusing on topics such as films, travel, and leisure activities.
@@
Table 2. Human evaluation results of phrase extraction for directly prompting GPT-4o vs. our pipeline.
Method $\text{R}_{\mathcal{P}}$ $\text{P}_{\mathcal{P}}$ $\text{F1}_{\mathcal{P}}$ $\Delta\text{F1}_{\mathcal{P}}$
Directly prompting GPT-4o 65.09 87.92 74.80 +00.00
Phrase extraction in pipeline 97.21 85.52 90.99 +16.19
      + Filtering rules 95.13 92.32 93.70 +18.90
Comparison with Human Annotation. To evaluate the privacy leakage classification step, we have 8 annotators classify 400 samples with half in Chinese and half in English, labeling them as either leak or non-leak. The guidelines to ensure the objectivity and consistency of the annotation results are provided in Appendix C. The results show an accuracy of 0.96, a recall score of 0.97, and a precision score of 0.94. These metrics demonstrate that the results of the first step closely align with the manual annotations.
@@
We also conduct a human evaluation for the privacy phrase extraction step, as shown in Table 2. Specifically, we manually extract privacy phrases for 200 samples to serve as the ground truth, resulting in 351 privacy phrases. Using the evaluation method described in Section 4.4, we found that the recall score for phrases directly extracted by GPT-4o with prompts (Zhang et al., 2024) was only 65.09%, meaning that roughly one-third of privacy phrases were missed. In comparison to this baseline, our pipeline, by incorporating category information as additional input, shows a slight decrease in precision but significantly increases the number of extracted phrases, boosting the recall score by 32.12%. As a result, the F1 score improves by 16.19%. Furthermore, applying filtering rules further enhances the F1 score by an additional 2.71%. These results also suggest that the phrase extraction step aligns well with manual annotations, highlighting its effectiveness and consistency.
To further evaluate the quality of the annotated privacy information, we randomly select 200 samples that contain 427 pieces of private information and ask annotators to evaluate the privacy information. Each information is scored as 1 if the generated privacy information accurately reflects the private details of the phrase and is semantically coherent; otherwise, it receives a score of 0. The average final score is 0.93, indicating high-quality and reliable annotation of the privacy information.
Table 3. Accuracy (%) of different baselines for privacy leakage classification.
Method
Model
English Chinese
Acc
Acc
ZG (Brown et al., 2020) Llama-3.2-1B $61.18_{\pm 0.49}^{*}$ $58.08_{\pm 0.49}^{*}$
Llama-3.2-3B $5.07_{\pm 0.48}^{*} 66$$6.97_{\pm 0.47}$
@@
Table 4. Performance of different privacy phrase extraction baselines from phrase-level metrics (%).
Method
Model
English Chinese
$\text{R}_{\mathcal{P}}$
$\text{P}_{\mathcal{P}}$
$\text{F1}_{\mathcal{P}}$
$\text{R}_{\mathcal{P}}$
$\text{P}_{\mathcal{P}}$
$\text{F1}_{\mathcal{P}}$
ZG (Brown et al., 2020) Llama-3.2-1B $29.67_{\pm 0.43}^{*}$ $27.21_{\pm 0.38}^{*}$ 28.39 $22.07_{\pm 0.40}^{*}$ $26.08_{\pm 0.43}^{*}$ 23.91
@@
4.4. Evaluation Metrics
Privacy Query-Level Metric for Leakage Classification Task. Given a user’s query, the objective of the privacy leakage classification task is to determine whether the query reveals any privacy of the user or user-related individuals. The automatically annotated data can be divided into either “leaked” or “non-leaked,” based on whether there are privacy phrases extracted. For such query-level classification task, we adopt accuracy to measure the performance of different local privacy detection methods.
Privacy Phrase-Level Metrics for Phrase Extraction Task. Given a private query, the objective of the privacy phrase extraction task is to extract phrases based on the context of the query that result in the user’s privacy leakage. For each data sample, we focus on both missed and incorrect extractions. For missed extractions, we define the recall score of sample $q$ as the ratio of the intersection between the predicted phrases and the ground truth phrases, normalized by the total number of ground truth phrases. For incorrect extractions, we define the precision score of $q$ as the ratio of the intersection between the predicted phrases and the ground truth phrases, normalized by the total number of predicted phrases. The overall recall and precision scores are calculated as the average of these scores across all samples.
Specifically, we let $\mathcal{Q}_{r}$ denote the set of user queries where the extracted phrase is non-empty in the ground truth and $\mathcal{Q}_{p}$ denote the set of user queries where the extracted phrase is non-empty in the prediction results. We let $\hat{\mathcal{P}}_{q}$ and $\mathcal{P}_{q}$ denote the set of predicted phrases and the ground truth of the query $q$, respectively. Then, the recall score $\text{R}_{\mathcal{P}}$ and precision score $\text{P}_{\mathcal{P}}$ is defined as
(1) $\text{R}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\mathcal{P}_{q}|},\ \ \text{P}_{\mathcal{P}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{q\in\mathcal{Q}_{p}}\frac{|\hat{\mathcal{P}}_{q}\cap\mathcal{P}_{q}|}{|\hat{\mathcal{P}}_{q}|},$
where $|\cdot|$ is the size of the set. In practice, the extracted phrases may refer to the same entity but have different expressions. To address this issue, our phrase matching strategy goes beyond simply checking whether two phrases are identical. It also considers two phrases a match if one is part of the other or if their Rouge-L score exceeds 0.5 (Huang et al., 2023b).
Privacy Information-Level Metrics for Information Summarization Task. Given a query with privacy phrases, the objective of the privacy information summarization task is to summarize specific information that is leaked from the query. In general, we use Rouge-L to evaluate the quality of generated privacy information. Similarly, we define the miss and incorrect extractions of the privacy information. We let $\hat{\mathcal{I}}_{q}$ and $\mathcal{I}_{q}$ denote the set of predicted information and the ground truth of the query $q$, respectively. For one privacy information ${I}_{j}\in{\mathcal{I}}_{q}$, the recall score is defined as $f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})=\max_{\hat{I}_{k}\in\hat{\mathcal{I}}_{q}}\text{{Rouge}}({I}_{j},\hat{I}_{k}),$ where Rouge is the function that calculates the Rouge-L score for two sentences. The recall score of ${\mathcal{I}}_{q}$ is defined as the average recall score of each piece of information it contains. For $\hat{I}_{j}\in\hat{\mathcal{I}}_{q}$, the precision score is defined as $f_{\text{P}}(\hat{I}_{j},\mathcal{I}_{q})=\max_{I_{k}\in\mathcal{I}_{q}}\text{{Rouge}}(\hat{I}_{j},I_{k}).$ The precision score of $\hat{\mathcal{I}}_{q}$ is defined as the average precision score of each piece of information it contains. Then we can define the recall score $\text{R}_{\mathcal{I}}$ and the precision score $\text{P}_{\mathcal{I}}$ for all queries as follow,
(2) $\text{R}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{r}|}\sum_{q\in\mathcal{Q}_{r}}\frac{1}{|\mathcal{I}_{q}|}{\sum_{I_{j}\in\mathcal{I}_{q}}f_{\text{R}}({I}_{j},\hat{\mathcal{I}}_{q})},$
$\text{P}_{\mathcal{I}}=\frac{1}{|\mathcal{Q}_{p}|}\sum_{q\in\mathcal{Q}_{p}}\frac{1}{|\hat{\mathcal{I}}_{q}|}{\sum_{\hat{I}_{j}\in\hat{\mathcal{I}}_{q}}f_{\text{P}}(\hat{I}_{j},{\mathcal{I}}_{q})}.$
Intuitively, $\text{R}_{\mathcal{I}}$ measures the completeness of the extracted privacy information, while $\text{P}_{\mathcal{I}}$ evaluates the accuracy of the information.
After getting the recall and precision score, we can also define the F1 score for both phrase-level and information-level as follows,
(3) $\text{F1}_{\mathcal{P}}=\frac{2\text{R}_{\mathcal{P}}\text{P}_{\mathcal{P}}}{\text{R}_{\mathcal{P}}+\text{P}_{\mathcal{P}}},\ \ \text{F1}_{\mathcal{I}}=\frac{2\text{R}_{\mathcal{I}}\text{P}_{\mathcal{I}}}{\text{R}_{\mathcal{I}}+\text{P}_{\mathcal{I}}}.$
5. Evaluation for Local Detection Methods
In this section, we evaluate various local privacy detection approaches based on the automatically annotated datasets.
5.1. Setups